I need help answering this question. I really do not understand.
If two numbers are opposite of each other, are their reciprocals also opposite of each other?
: I need help answering this question. I really do not understand.
If two numbers are opposite of each other, are their reciprocals also opposite of each other?

Answer by solver91311(1850)

For any real number a, -a is its opposite, meaning a + (-a) = 0. The reciprocal of a is 1/a, and the reciprocal of -a is 1/(-a), but 1/(-a)=(-1)/a=-1/a and 1/a+(-1/a)=0.